the larger of two numbers is 1 less than 3 times the smaller number. If twice the smaller number is increased by the larger number, the result is 18 more than the larger number. find both numbers
Answer by fractalier(6550)

Call the larger x and the smaller number y. Then we have
x = 3y - 1
and
2y + x = 18 + x
from which we get
2y = 18
y = 9
Plugging that in above we get
x = 3(9) - 1 = 26